Let f (x) = sqr symbol x + 6 and g(x) = x – 8

(a) Find (f/g) (-2) show work
(b) Find the domain of the quotient function f/g. Explain

Answer by Aldorozos(172)

Divide both equation
Sqrt x+6/x-8
Replace x with -2
4/-10 = -2/5= - .4
Domain: the denominator can't be zero
Sqrt of a number can't be negative which means under the sqrt can't be negative
X-8=8 therefor x can't be eight since 8 can make the denominator as zero
Sqrt x+ 6 has to be greater than zero. X+6= 0 therefore x can be -6 or any number greater than 6 except 8